OpenStudy (anonymous):

What is the inverse Laplace of (5-s)/(s^2+6s+9)?

OpenStudy (abb0t):

try and fctor your s^2+6s+9

OpenStudy (anonymous):

(5-s)/(s+3)^2

zepdrix (zepdrix):

If you split up your fractions, you can do these pieces separately.\[\large \frac{5}{(s+3)^2}-\frac{s}{(s+3)^2}\] Hmm I can explain the first piece. I'm trying to remember how to do the second one though. heh.

OpenStudy (abb0t):

note your denom \[(\frac{ 1 }{ s+3 })(\frac{ 1 }{ s+3 })\] which is familiar

OpenStudy (abb0t):

But, the bottom is 3, so you should miltiply by a number to get it into the form to get an exponential

OpenStudy (abb0t):

Im also trying to remember how to work this.

zepdrix (zepdrix):

\[\huge \mathcal L\left[t\cdot f(t)\right]=(-1)\frac{d}{ds}F(s)\]Does this formula look familiar yury? :o

OpenStudy (anonymous):

maybe partial fraction

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

We can use that on the first part, but in reverse.

zepdrix (zepdrix):

We're basically integrating the term, and attaching a d/ds to it. :D

zepdrix (zepdrix):

\[\large \frac{5}{(s+3)^2} \qquad \rightarrow \qquad (-1)\frac{d}{ds}\frac{-5}{(s+3)}\]

zepdrix (zepdrix):

Understand what I did there? :o I integrated, but attached a derivative operator to it at the same time, so it hasn't affected the value of the fraction.

OpenStudy (anonymous):

the question starts with y"+6y'+9y=0 y(0)=-1 y'(0)=6

OpenStudy (anonymous):

the answer in the book is -e^(-3t)+3te^(-3t)

zepdrix (zepdrix):

Yes, see the second term? That's where I was headed with the steps I was showing you.

zepdrix (zepdrix):

err wait, we would end up with a 5 instead of a 3 coefficient, so yah you musta run into a tiny mistake somewhere earlier on.

zepdrix (zepdrix):

Hmm I don't think that 5 should be there, check to see if you forgot to distribute the 6 in the second term.